OpenStudy (anonymous):

I was told this question was between the 4th and 1st quadrant...i know the answer i just want to know how i know which quadrant an equation takes place in (equation posted soon)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}x^3/\sqrt{x^2+25}dx x=5\tan\]